The Ksp of Mg(OH)2 is 8.9 10-12 at 25*C. If the pH of solution is adjusted to 9.0 . How much Mg+2 ion will be precipitated as Mg(OH)2 from a 0.1 M MgCl2 solution at 25*C ? Assume that MgCl2 is completely dissociated.
​ a.  0.011
​b.  0.089
c.  0.11
d.  0.89
Please explain in detail............. 

Grade:12th pass

1 Answers

Vikas TU
3 years ago
Ksp[Mg(OH)2]=8.9×10^−2 
pH=9
[H+]=10^−9
[OH−]=10^−5
Ksp=[Mg2+][OH−]^2
8.9×10^−12=[Mg2+][10^−5]^2
[Mg2+]=8.9×10^−2=0.089
Conc. Of Mg2+ precipitated =0.1−0.089
=0.011mol/L.t